OpenStudy (anonymous):

a rectangle is 3cm longer than it is wide. the diagonal is 15cm. find the dimensions of the rectangle

OpenStudy (anonymous):

I thought, it was a bit odd. :) Solve simultaneous equations: Width + 3 = Height Width^2 + Height ^2 = 15^2

OpenStudy (anonymous):

let the width be =x cm then length is = (x+3) cm Withe the diagonal, length and width form a right angled triangle whose hypotenuse is the diagonal Using Pythagoras theorem (15)^2 = (x)^2 + (x+3)^2 225 = x^2 + x^2 + 6x + 9 225 = 2x^2 + 6x + 9 or 2x^2 + 6x - 216 = 0 Factorising the quadratic equation 2x^2 + 6x - 216 = 2(x^2 + 3x - 108) = 2(x^2 + 12x - 9x - 108) = 2[x(x + 12) -9(x + 12)] = 2[(x - 9)(x + 12)] so 2(x-9)(x+12) = 0 => either x-9 = 0 or x+12 = 0 if x-9 = 0 ==> x = 9 if x+12 = 0 => x = -12 since length can never be negative, so the usable value is x = 9 so width = 9 cm length = 9+3 = 12 cm